f(x) = √(tanx) + √(sinx) dx
1
Expert's answer
2020-05-31T15:11:54-0400
"f(x)=\sqrt{\tan(x)}+\sqrt{\sin(x)}"

"{df \over dx}={1 \over 2\sqrt{\tan(x)}}({1 \over \cos^2x})+{1 \over 2\sqrt{\sin(x)}}(\cos(x))"